OpenStudy (anonymous):

Let log_bA=3 log_bC=2 log_bD=5 then what is the value of log_b D^2/ C^3 A

OpenStudy (anonymous):

\[\log_b \frac{ D^2 }{ C^3A }\]

OpenStudy (anonymous):

thats what i am looking for the value for ^^^^^

OpenStudy (mathmate):

Use the laws of logarthms: \[log(ab)=log(a)+log(b)\] \[log(a^b)=b\ log(a)\] to simplify the expression and substitute the given values.

OpenStudy (anonymous):

mmmhhh ok then

geerky42 (geerky42):

He forgot one more: \(\log\left(\dfrac{a}{b}\right) = \log(a) - \log(b)\) You would need to apply that first. Can you do it?

OpenStudy (anonymous):

mmmhhh i am attempting right now .... lol would you mind checking if i got it when im done ??? @geerky42

geerky42 (geerky42):

Sure

OpenStudy (mathmate):

@geerky42 yep, good catch!

OpenStudy (anonymous):

lol soooo i did all of the properties and something went wrong my answer was not one of the answers provided .... heres what i did \[\log_b \frac{ D^2 }{ C^3 A }\] \[\log_bD^2-\log_bC^3A\] \[2\log_bD-\log_bC^3+\log_bA\] \[2\log_bD-3\log_bC+\log_bA\] from here i simply added the numbers ... 2(5) - 3(2) + 3 10-6+3 = 7

OpenStudy (anonymous):

@geerky42 @mathmate okkk technical difficulties .... -_- lol

OpenStudy (mathmate):

I think the sign for log(A) should be negative...

OpenStudy (mathmate):

... unless you add parentheses around (3log(C)+log(A))

geerky42 (geerky42):

+logbA is error here.\[-\log_bC^3A = -(\log_bC^3A) = -(\log_bC^3+\log_bA) = -\log_bC^3 -\log_bA\]

OpenStudy (anonymous):

ohhhh didnt apply the negativeeeee ohhh ok

OpenStudy (anonymous):

ohh then its one !

geerky42 (geerky42):

Right!

OpenStudy (anonymous):

yay thats actually an answer given thankkkk you n_n !!@geerky42 eheheh i did it n_n

OpenStudy (anonymous):

thank you n_n @geerky42 @mathmate

OpenStudy (mathmate):

You're welcome! :)